3 ( log 5 - log 3 ) - ( log 5 - 2 log 6 ) = 2 - log n , find n.

Answer by dkppathak(439)

3 ( log 5 - log 3 ) - ( log 5 - 2 log 6 ) = 2 - log n , find n.
3log 5/3 -lg5/36 + log n =2
log 125/27 -log 5/36 +log n =log100
log 125x36/5x27 +log n =log100
log 100/3 +log n =log100
log100n/3=log100
100n/3=100
n=100x3/100 =3
ANSWER N = 3